# Self-Specialization: Uncovering Latent Expertise within Large Language Models

### v1

@@ Appendix B Details of Experiments
In Table 3, we show the prompts used for our self-specialization. For instruction generation, we leverage the prompt designed in self-instruct Wang et al. (2022a) with minimal change to make it suit to specialization. In particular, we ask a model for instructions about a targeted domain, and force it to generate input together with the instruction, unlike in Wang et al. (2022a) that generates those separately. In addition, we avoid using the specific requirement in the prompt that asks to cover diverse topics, such as (quoting Wang et al. (2022a)) “daily routines, travel and tourism health and wellness, cooking and recipes, personal finance, environmental issues, history and historical events, literature and literary analysis, politics and current events, psychology and mental health, art and design, mathematics and problem-solving, physics and astronomy, biology and life sciences, chemistry and materials science, computer science and programming, engineering and technology, robotics and artificial intelligence, economics and business management, philosophy and ethics, and more”. For response generation, we use a simple prompt to let a model answer with a target domain in mind. Both prompts can be further enhanced and optimized for better self-specialization performance in future work.
@@
Regarding our evaluations, we use prompt templates that were designed and used to optimize each Alpaca (Taori et al., 2023) and Dromedary (Sun et al., 2023), but no specific template for base models, as they were not optimized for it during pre-training. Ours employs a simple Alpaca template for training and evaluation. We leverage publicly available delta weights that are supposed to be attached to LLaMA (Touvron et al., 2023) for Dromedary, and use the ones reproduced for Alpaca in our work.
@@
We use three seed demonstrations in-context, which are randomly sampled from our initial seeds, and sampling with top-p being 0.98 and temperature being 1.0 during instruction generation. For response generation, we use no demonstrations in-context since there is a high chance that the generated instruction task and the sampled one do not match well. We believe further exploration of this aspect would be valuable in future work. For fine-tuning, we use a batch size of 32, a learning rate of 3e-4, and epochs of 3. Low-rank adaptation (LoRA) (Hu et al., 2022; Dettmers et al., 2023) is applied to all modules and all layers with a rank of 8, and an alpha of 16.

### v2

@@ Appendix B Details of Experiments
In Table 10, we show the prompts used for our self-specialization. For instruction generation, we leverage the prompt designed in self-instruct Wang et al. (2022a) with minimal change to make it suit to specialization. In particular, we ask a model for instructions about a targeted domain, and force it to generate input together with the instruction, unlike in Wang et al. (2022a) that generates those separately. In addition, we avoid using the specific requirement in the prompt that asks to cover diverse topics, such as (quoting Wang et al. (2022a)) “daily routines, travel and tourism health and wellness, cooking and recipes, personal finance, environmental issues, history and historical events, literature and literary analysis, politics and current events, psychology and mental health, art and design, mathematics and problem-solving, physics and astronomy, biology and life sciences, chemistry and materials science, computer science and programming, engineering and technology, robotics and artificial intelligence, economics and business management, philosophy and ethics, and more". For response generation, we use a simple prompt to let a model answer with a target domain in mind. Both prompts can be further enhanced and optimized for better self-specialization performance in future work.
@@
Regarding our evaluations, we use prompt templates that were designed and used to optimize each Alpaca (Taori et al., 2023) and Dromedary (Sun et al., 2023), but no specific template for base models, as they were not optimized for it during pre-training. Ours employs a simple Alpaca template for training and evaluation. We leverage publicly available delta weights that are supposed to be attached to LLaMA (Touvron et al., 2023a) for Dromedary, and use the ones reproduced for Alpaca in our work.
@@
We use three seed demonstrations in-context, which are randomly sampled from our initial seeds, and sampling with top-p being 0.98 and temperature being 1.0 during instruction generation. For response generation, we use no demonstrations in-context since there is a high chance that the generated instruction task and the sampled one do not match well. We believe further exploration of this aspect would be valuable in future work. For fine-tuning, we use a batch size of 32, a learning rate of 3e-4, and epochs of 3. Low-rank adaptation (LoRA) (Hu et al., 2022; Dettmers et al., 2023) is applied to all modules and all layers with a rank of 8, and an alpha of 16. While we report single-run results considering low-data settings where automatic hyperparameter tuning might be infeasible, we also report worst, average, and best across different k-shot configurations for each dataset to address the concern of sensitivity (Appendix C.2) in Table 5.
-Score / Rouge-L Task Dataset Base Self-Specialized QA BioASQ-Factoid 51.96 / 51.81 57.61 / 57.48 BioASQ-List 35.09 / 30.40 42.17 / 36.24 BioASQ-Yesno 8.80 / 8.80 95.20 / 95.20 PubMedQA 31.69 / 24.56 31.31 / 24.77 NER AnatEM 6.59 / 6.07 21.25 / 19.24 BioNLP13CG 26.03 / 22.53 41.16 / 35.07 NCBI 17.99 / 16.60 46.54 / 41.55 RE DDI 49.38 / 49.38 53.40 / 53.40 SA Medical Drugs 11.40 / 11.40 32.80 / 32.80 DC HoC 62.84 / 62.84 62.65 / 62.65 Average 30.18 / 28.44 48.41 / 45.84
Task Dataset Base Self-Specialized SOTA QA BioASQ-Factoid 51.96 57.61 49.51 BioASQ-List 47.57 46.99 35.59 BioASQ-Yesno 21.20 95.20 68.25 PubMedQA 31.69 31.31 29.58 NER AnatEM 9.63 21.25 84.61 BioNLP13CG 26.03 41.16 65.09 NCBI 27.88 46.54 80.91 RE DDI 51.00 53.40 89.35 SA Medical Drugs 35.00 65.80 47.37 DC HoC 62.84 62.65 82.53 Average 36.48 52.19 63.23
Biomedicine Dataset Base Self-Specialized AdaptLLM QA BioASQ-Factoid 39.21 52.17 51.79 BioASQ-List 32.45 43.99 49.74 BioASQ-Yesno 66.00 88.40 93.80 PubMedQA 23.59 31.04 24.06 NER AnatEM 1.20 20.93 9.81 BioNLP13CG 22.16 31.46 18.58 NCBI 37.91 43.28 20.50 RE DDI 47.6 53.80 51.00 SA Medical Drugs 42.80 24.20 13.60 DC HoC 10.60 50.87 8.33 Average 32.35 44.01 34.12 Finance Dataset Base Self-Specialized AdaptLLM SUM EDT-Summarization 13.87 23.56 35.95 QA ConvFinQA 28.77 20.88 12.10 InsuranceQA 9.96 24.36 26.11 NER Fin3 6.80 43.87 13.32 FiNER_139 44.34 35.63 11.10 RE KPI-EDGER 49.46 63.90 34.53 SA EarningsCall 49.03 47.74 47.30 Financial_Phrasebank 29.20 68.80 34.80 FIQA-SA 61.54 70.09 69.66 CLS Gold Commodity News 38.42 61.20 71.61 Average 33.14 46.00 35.65
Models # Labeled Data # Synthetic Data MMLU Score Humpback-65B (General) 3200 500K 59.0 Ours-7B (Specialized) 100 5K 64.0
Appendix C Additioanl Results & Discussion
C.1 Qualitative Analyses
While our study primarily focuses on the biomedical and finance domain, the applicability and effectiveness of self-specialization in another specialized domain whose knowledge is relatively limited, such as sports, remain an open avenue for exploration. As an initial effort, we present a case study of a self-specialized model on sports in Table 13 & 14, along with the visualization of generated data in Figure 8. We hope that this could offer insights into the versatility of self-specialization, although the model is not yet perfect, and thorough evaluations are required in future work. Different domains inherently pose unique requirements and nuances, and understanding how self-specialization adapts to these variations is a valuable direction for future work.
C.2 On the Sensitivity of Prompting
In Table 2, we observe the decreased performances with increased demonstrations in certain cases such as BioASQ and Medical Drugs. We conjecture this can be attributed to the model’s sensitivity (Zhao et al., 2021) or interference among demonstrations (Chen et al., 2023) under in-context learning (ICL). In fact, it can even be noticed in the original GPT-3 paper (Brown et al., 2020) that additional demonstrations do not always lead to better performance and can indeed sometimes result in a notable decrease, demonstrating an inherent challenge in ICL. Taking the worst, average, and the best across different k-shot (0, 1, 5) configurations for each dataset to address the concern of sensitivity, we still notice the significant gaps between our self-specialization and the base model, presented in Table 5.
C.3 On Evaluation Designs
In our study, as described in Section 4, we treat all tasks as a unified text generation problem, aiming to assess the realistic capabilities of following instructions, consistent with established practices in biomedical instruction tuning literature (Parmar et al., 2022). As briefly discussed in Section 5.1, we observe that in Table 2, the base model’s performance on BioASQ-Yesno is very low (below random), often failing to follow instructions and generating text that is not confined to the label space. We therefore treat this dataset as an outlier and exclude it from our average calculations. Even after removing this outlier, self-specialization still has substantial gains over the base model: 25.58 to 31.22 (0-shot), 28.42 to 36.55 (1-shot), and 32.55 to 43.21 (5-shot). However, we believe that our current evaluation is fairer and preferable, because in a realistic scenario where a user prompts a model to solve a certain task (e.g., classification) without the assumption about a task type, and gets a totally wrong response out of the label space, evaluating such a response as correct would not make sense.
The primary objective of our work is to enhance the base model’s domain-specific capabilities through self-specialization, a process inherently different from conventional fine-tuning approaches. Although the process utilizes LoRA for specialization, it is important to note that our approach fundamentally relies on synthetic data generated by the model itself. This unique aspect sets our method apart, as it effectively starts from scratch, focusing on self-generated, domain-specific instructional data for low-data scenarios. Finally, the base model and the base model improved through our Self-Specialization (using synthetic self-generated data) are compared fairly in the same zero-shot/few-shot setting.
C.4 Comparison with State-of-the-art
In Table 7, we present the performances of a state-of-the-art instruction-tuned model (Parmar et al., 2022) in a biomedical domain, for a reference point. It is important to clarify that our comparison should not be considered direct. The SOTA model, unlike ours which relies on a few seed samples, is fine-tuned on a vast corpus of human-annotated data (140K), and differences in test set splits may exist. For the base MPT and self-specialized models, the maximum performances by using up to 5 samples are presented.
Despite significant improvements over its base model, our self-specialized model remains behind SOTA benchmarks, which is not surprising due to the nature of our method that is not supervised, unlike the SOTA model. While expected, this possibly implies the practical utility of our approach may be limited yet in certain scenarios. From the table, we especially note the substantial gap in Named Entity Recognition (NER) tasks. This gap can be attributed to the SOTA model’s training on a large and diverse set of NER datasets (i.e., 80K samples). This suggests ample opportunity for further exploration and enhancement in this area.
Additionally, we provide comparisons with concurrent works (Cheng et al., 2024; Li et al., 2024) in Table 8 and 9. To make distinctions, those works are, in principle, oriented toward different scenarios from ours. Cheng et al. (2024) leverages a large specialized corpus to transform them into reading comprehension texts using pre-defined templates, which are then used for training LLMs. Differently, our work does not necessitate such a huge amount of corpus that requires 768 GPU hours for training; ours takes only a few GPU hours. Moreover, we do not confine our method to specific tasks, aiming for cross-task generalization as shown in Table 8, unlike them focusing on constrained reading comprehension format.
Li et al. (2024) explores the augmentation of instructional data through the use of a few thousand labeled seed data to initialize their generation pipeline relying on a web corpus. Unlike them assuming huge amounts of seed/unlabeled data, our method requires only a handful of seeds (e.g., <100). Furthermore, their focus is on general instructions, which have shown only marginal effects in specialized domains in our preliminary results (Section 2). Given the infeasibility of direct comparison, we reference their MMLU result from the paper to contextualize and we self-specialize Gemma-7B (Team et al., 2024) on MMLU (Hendrycks et al., 2021). While acknowledging their values in different scenarios from ours, we believe this distinction provides an insight into the unique value and impact of our work in enhancing model performance along with efficiency in targeted domains.